What is the molar solubility of manganese (II) carbonate in a solution of 0.250M sodium carbonate? Would the molar solubility be greater or smaller is no sodium carbonate was added? Ksp=2.34*10**-11

Please explain!
Answer #1

You have the reaction:

MnCO3 = Mn + 2 + CO3-2

You have the expression of Ksp:

Ksp = [Mn + 2] * [CO3-2] = X * (0.25 + X) = 2.34x10 ^ -11

It clears X = 9.4x10 ^ -11 M

The solubility will be lower, since there is a common ion effect CO3-.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.